OpenStudy (anonymous):

Find the angle between the vectors for u and v given below. u=<4,3> v=<-2,4>

OpenStudy (anonymous):

\[u \cdot v=||u||*||v||\cos(\theta)\] So: \[\arccos(\frac{u \cdot v}{||u||*||v||})=\theta\] \[u \cdot v=(4)(-2)+(3)(4)=-8+12=4\] \[||u||=\sqrt{16+9}=5\] \[||v||=\sqrt{4+16}=\sqrt{20}\] So: \[\theta=\arccos(\frac{4}{5\sqrt{20}})=79.7 \] degrees

OpenStudy (anonymous):

Thank you very much. I had it right I was just taking cos not arccos. thank you very much

OpenStudy (anonymous):

No problem :P